OpenStudy (anonymous):

Hey, can someone please explain how to differentiate 3rootx times e^x plz?

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] with \[f(x)=\sqrt{x}, f'(x)=\frac{1}{2\sqrt{x}}, g(x)=e^x, g'(x)=e^x\]

OpenStudy (anonymous):

my teacher told us that the answer was (1+3x)e^x/3(3rootx^2) ... Not sure how he gets this?